TPC

ALTRO Settings

I toyed with Yuri's TPC simulation curve and I show you just one result:


Blue line (before filter):
--------------------------

Using Yuri's curve and multiplying it to
suit my needs I added two hits together.

The first peak simulates a highly ionizing track.
Note how high the peak ADC is.

Second peak simulates a MIP (obtained by
dividing the first by 10, moving 11 timebins
to the right and adding up).



Red line (after filter):
-----------------------

You can see the baseline restoration with a minimal
loss of signal!

Not bad.

BTW, the ALTRO parameters were:

K1 12807
K2 50251
K3 64797
L1 65145
L2 54317
L3 8537

       -- Tonko

 

FCF

 Please attach any information about the Fast Cluster Finder (FCF) here.

Cluster flags

TPC forum post 425:

I am sending a short description of the
TPX cluster finder flags. This is more an Offline issue
but I guess this is a better group since it is a bit
technical.

1) The flags are obtained in the same way other
variables such as timebin and pad are obtained.
I understand that this is not exported to StHit *shrug*.

See i.e. StRoot/RTS/src/RTS_EXAMPLE/rts_example.C

2) They are defined in:

StRoot/RTS/src/DAQ_TPX/tpxFCF.h

The ones pertaining to Offline are:

FCF_ONEPAD           This cluster only had 1 pad.
Generally, this cluster should be ignored unless
you are interested in the prompt hits where this
might be valid. The pad resolution is poor, naturally.

FCF_MERGED           This is a dedconvoluted cluster.
The position and charge have far larger errors
than normal clusters.

FCF_BIG_CHARGE    The charge was larger than 0x7FFF so
the charge precision is lost. The value is OK but the precision is
1024 ADC counts. Good for tracking, not good for dE/dx.

FCF_BROKEN_EDGE       This is the famous row8 cluster.
Flag will disappear from valid clusters
once I have the afterburner running.

FCF_DEAD_EDGE          Garbage and should be IGNORED!
This cluster touches either a bad pad or an end
of row or is somehow suspect. I need this flag for internal debugging
but the users should IGNORE those clusters!

        -- Tonko


TPC forum post 426:

I commited the "padrow 8" afterburner to the DAQ_TPX
CVS directory. If all runs well, you should see no more
peaks on padrow 8. The afteburner runs during the
DAQ_READER unpacking.

However, please pay attention to the cluster finder flags
which I mentioned in an email ago. Specifically:

"FCF_DEAD_EDGE          Garbage and should be IGNORED!
This cluster touches either a bad pad or an end
of row or is somehow suspect. I need this flag for internal debugging
but the users should IGNORE those clusters!"

         -- Tonko

Field Cage Shorts

This page is for information regarding shorts or current anomalies in the TPC field cages.

Excess current seen in 2006

The attached powerpoint file from Blair has plots of the excess current seen in the IFC East for 2006.

Modeled distortions

Modeling the Distortion

Using StMagUtilities, Jim Thomas and I were able to compare models of the distortions caused by shorts at specific rings in the IFC with the laser data. First, I'll have to say that I was wrong from my Observed laser distortions: the distortion to laser tracks does not have the largest slope at the point where the short is. Instead, it has a maximum at that point! The reason is that the z-component of the electric field due to the distortion (withouth compensating resistor) changes signs at the location of the short. So ExB also changes directions, and the TPC hits are distorted in one rPhi direction on the endcap side of the short, and in the opposite rPhi direction closer to the central membrane.

This can be seen in the following plot, where I again show the distortion to laser tracks at a radius of 60cm (approximately the first TPC padrow) versus Z in the east TPC using distored run 7076029 minus undistorted run 7061100 as red data points. Overlayed are curves for the same measure from models of a half-resistor short (actually, 1.14 MOhm short as determined by the excess current of ~240+/-10nA [a full 2MOhm short equates to 420nA difference]) located at rings 9.5, 10.5, ..., 169.5, 179.5 (there are only 182 rings).

[note: earlier plots I have shown included laser data at Z = -55cm, but I've found that the laser tracks there weren't of sufficient quality to use; I've also tried to mask off places where lasers cross over each other]

The above plot points towards a short which is located somewhere among rings 165-180 (Z < -190cm). As the previous years' shorted rings were rings 169 and 170 ( = short at ring 169.5), it seems highly likely that the present short is in the same place. More detail can be seen by looking at the actual laser hits. The first listed attached file shows the laser hits as a function of radius for lasers at several locations in Z. The dark blue line is a simple second order polynomial fit I used to obtain the magnitude of the distortion at radius 60cm, which I used in the above plot. The magenta line is the model of the half-resistor short at ring 169.5, and the light blue line is the same for ring 179.5 (the bottom two curves on the above plot). Either curve seems to match the radial dependence fairly well.

Further refinement can be achieved by modeling the exact resistor chain. We have a permanent short at ring 169.5 (rings 169 and 170 have been tied together), and have replaced the two 2MOhm resistors between 168-169 and 170-171 with two 3MOhm resistors (see the attached photo of the repair, with arrows pointing to candidate locations for shorts via drops of silver epoxy). So it is more likely that we have a 1.14MOhm short on one of these two 3 MOhm resistors. The three curves in this next plot are:

  • red: 1.14MOhm short on a 3MOhm resistor at 168.5, full short at 169.5, 3MOhm resistor at 170.5
  • green: 1.14MOhm short on a 2MOhm resistor at 169.5, normal 2.0MOhm resistors at 168.5 and 170.5
  • blue: 3MOhm resistor at 168.5, full short at 169.5, 1.14MOhm short on a3MOhm resistor at 170.5
The difference between the curves does not come from the fact that we have treated the 3MOhm resistors properly (that difference is less than 5 microns in distortions, and only in a small region near the short! shown here), but rather from the movement of the effective short from location 168.5 to 169.5 to 170.5. From a visual inspection (not a fit), it appears that the blue curve is the best match to the data:


We can also take a look at the data with the resistor in. Here is the same plot as before with a 1.14MOhm short at the same locations, but with an additional compensating resistor of 1.0MOhm. The fact that all the data points are below zero points again towards a short near the very end of the resistor chain, preferring a location of perhaps 177.5 over shorts near ring 170. These plots do not include the use of the 3MOhm resistors, but that difference is below the resolution presented here.

Zoom in with finer granularity between rings (every other ring short shown):

The second listed attached file shows the laser hits as a function of radius for lasers at several locations in Z for the case of the resistor in, again with magenta and blue curves for the model with shorts at ring 169.5 and 179.5 respectively.


Applying the Correction

I tried running reconstruction on the lasers using the distortion corrections for the 1.14MOhm short at three locations: 170.5 and 171.5 (two possible spots indicated in Alexei's repair photo), and 175.5 (closer to what the with-resistor data pointed to). The results are in the following plots. The conclusion is that the 175.5 location seems to do pretty well at correcting the data, slightly better than the 170.5 and 171.5 locations, for both with and without compensating resistor. For this reason (the laser data), we will proceed with FastOffline using a short at 175.5, even though we have no strong reasons outside the laser data to suspect that the short is anywhere other than the rings 168-172 area where the fix was made.

170.5171.5175.5


Gene Van Buren
gene@bnl.gov

Choosing an external resistance for an IFC short

It is clear that when a field cage short is close to the endcap, it is best to add an external resistance to compensate for the amount of the missing resistance from the short as that would restore the current along the length of the resistor chain and only disrupt the potential at the last (outermost) rings instead of along the full length. A short near the central membrane would benefit less from this as restoring the proper current does restore the potential drop between each ring, but leaves almost all rings at a potential offset from the intended potential, essentially tilting the E field over nearly the whole volume.

But the question then comes as to whether we can decide on the best external resistance for minimizing the distortion, to align with the principle that the best distortion we can choose is the one which requires the least correction, in case we're not quite correcting it accurately. To answer that, the distortion modeling was run with a variety of locations for an IFC west 2 MOhm single-resistor short, and a variety of external resistances. The code to run this modeling has been attached as a tar file to this Drupal page in case there is interest to re-run it (e.g. for an OFC short).

Here are the results:
  • Left: Surface plot showing the mean r-φ distortion we would get at the first iTPC pad row (radius = 55 cm) averaged over all active west-side z as a function of where the short is located ("ring of short") and how much external resistance is added.
  • Right: Same but drawn as contours. The curve of interest to follow, that leads to <distortion>=0, is the one that starts near (ring,external resistance) = (110,0.0) and ends near (180,2.0), as indicated by the red markers. If the short is at ring 160.5, for example, then this curve indicates an external resistance of ~1.05 MΩ minimizes the distortions averaged over all active west-side z.



However, it may be more important to restrict the z range included in the distortion average, as most tracks of interest do not cross the inner pad rows at high z...
  • Left: Similar surface plot, but restricting the average to 0 < z < 100 cm.
  • Right: The curve of interest to follow in this contour plot, that leads to <distortion>=0, is the one that starts near (100,0.0) and ends near (180,2.0), as indicated by the red markers. If the short is at ring 160.5, then this curve indicates that an external resistance of ~1.25 MΩ minimizes the distortions averaged over 0 < z <100 cm.




Some additional observations:
  • The curve of interest should always end close to (180.5,2.0), as that approaches the condition where the external resistance is no different than an internal resistance at the very end of the chain.
  • The above is a simplification of what area should be integrated, as tracks with η ≠ 0 cross a variety of z at various radii, complicating the impact on their reconstruction. A track-by-track analysis of impact would be more meaningful, but a lot more work! The modeling shown here can serve as a rough guide to the best external resistance to use, but should not be taken as definitive for all physics.
  • It is interesting to note that the model implies that a negative external resistance would help minimize the <distortion> when the short is closer to the central membrane (ring 0). A way to think of this is like having a short at both ends, such that the potentials are too high near the central membrane, and then too low near the endcap, so that the E field tilts one way in the region near the central membrane, isn't tilted at half the drift length, and then tilts the other way near the endcap, resulting in opposing distortions for electrons which drift the full length that serve to cancel each other. This could in principle be achieved by reducing the overall resistance of the Resistor box at the end of the Field Cage chain. The STAR TPC has (as of this documentation) had no persistent shorts near the central membrane that would warrant this approach.

-Gene

OFC West possible distortion

There remains a possible distortion due to a potential short in the OFC west as well. We see a bimodal pattern of 0 or 80 excess nanoAmps coming out of the OFC West field cage resistor chain (it has been there since the start of the 2005 run). That corresponds to a 0.38 MOhm short (420nA = 2 MOhms). The corresponding distortion depends on the location of the electrical short. The plot shown here is the distortion in azimuth (or rPhi) at the outermost TPC padrow near the sector boundaries (r=195 cm, the pads are closer to the OFC near the sector boundaries) due to such a short between different possible field cage rings:

In terms of momentum distortion, a 1mm distortion at the outermost padrows would cause a sagitta bias of perhaps about 0.5mm for global tracks (and even less for primary tracks), corresponding to an error in pt in full field data of approximately 0.006 * pt [GeV/c] (or 0.6% per GeV/c of pt). This is certainly at the level where it is worthwhile to try to fix the distortion if we can figure out where the short is. It is also at the level where we should be able to see with the lasers perhaps to within 50cm where the short is.

Just as a further point of reference, the plot for radius = 189cm, corresponding to the radius of the outermost padrow in the middle of a sector (its furthest point from the OFC) can be found here.

This possible distortion remains uninvestigated at this time.


Gene Van Buren
gene@bnl.gov

Other FC short distortion measurements

I considered the possibility that other measurements might help isolate the location of the short in the TPC. So, using the Modeled distortions, I modeled the effects of adding even more compensating resistance to the end of the IFC east resistor chain. Below are the results for shorts located at ring 165.5 (between rings 165 and 166), 167.5, 169.5, ..., 179.5 as indicated for the colored curves. All plots use a 1.0cm range on the vertical scale so that they can be easily compared. I had hoped that one resistance choice or another would cause more separation between the curves, giving better resolving power between different short locations. But this dependence is small, and actually seems to decrease a little with increased compensating resistance. Remember also that the last laser is at Z of about -173cm.

Also perhaps worth noting is that the 1.0MOhm compensating resistor probably helps reduce the distortions even more than an exact 1.14MOhm would. The latter causes the distortions not to change between the short location and the central membrane, but the former actually causes the distortion to re-correct the damage done between the short location and the endcap!

Compensating
resistance
[MOhms]
Distortion on first padrow vs. Z
0.0
1.0
1.14
2.0
4.0
20.0


Gene Van Buren
gene@bnl.gov

Observed laser distortions

First look

The first listed attached file was my initial look at distortions to laser tracks in the TPC (not that it is upside down from subsequent plots as I accidentally took the non-distorted minus the distorted here). These are radial tracks, and the plots are off the difference between run 7068057 (with excess current) and 7061100 (without excess current). Also, run 7068057 was taken with collisions ongoing, so there is some SpaceCharge effect as well (7061100 was taken without beam). I believe this explains the rotation of the tracks at low positive z (west side).

Second look: dedicated runs

On March 17, we took a couple laser runs without and with a compensating resistor to get the IFC east current at least approximately correct. I plotted 1/p of laser tracks and took the profile. Straight tracks give very low curvature = low 1/p. The distortion brings up the curvature, as can be seen in the IFC east without resistor. The same plot also shows large error bars for the negatively charged laser tracks because there aren't many: the curvature tends to bring them positive. The IFC west shows the appropriately low level without any distortion, but the timing on the west lasers was wrong, so they are not reconstructed where they are supposed to be in Z. I am uncertain whether this bears any relation to the odd behavior of the first laser on the west side (showing up here at Z of about +67cm). The IFC East plot for the no-compensating-resistor run also shows that 1/p begins to drop somewhere around Z of about -100cm. The short would be located where the largest slope occurs in this plot (because the distortion to tracks is an integral of the distortion in the field, and the short is where the field distortions are largest), but the data isn't strong enough to pin this down very well. The negative tracks indicate a short between the lasers at -145cm and -115cm. But the seemingly better quality positive tracks are less definitive on a location as the slope appears to get stronger at more negative Z, implying a short which is at Z beyond -170cm.

No compensating resistor (run 7076029):

With 1 MOhm compensating resistor (run 7076032), which brings IFC east current to correct value within ~20 nAmps, according to 10:27am 2006-03-17 entry in Electronic ShiftLog:

Again, I looked at the distortion as seen be comparing TPC hits from the distorted runs to those from an undistorted run as I did at the beginning of this page (but this time taking distorted minus undistorted). For the undistored, I again have only run 7061100 to work with as a reference. The plots for each Z are in the second listed attached file below, where the top 6 plots are for the no-compensating-resistor run (7076029), and the bottom 6 are the same ones with the resistor in. I also put on the plots the value of the difference from a simple fit (meant only to extract an approximate magnitude) at a radius of 60cm (approximately the first TPC pad row). Those values are also presented in the following plot as a function of Z, confirming the improvement of the distortion with the resistor in place.

These plots seem to point at a short which is occurring somewhere between the lasers at Z = -145 and -115cm.


Gene Van Buren
gene@bnl.gov

Resistor box at the end of the Field Cage chain

 After ring 181, the potentials are determined by a box of resistors which sit outside the TPC. This is well documented, but at the time of this writing is not complete. This was particularly relevant during Run 9 when an electrical short developed inside the TPC between rings 181 and 182 of the outer field cage on the west end (OFCW). Shown here is a plot of the resistors:

 

Note that the ammeter is essentially a short to ground, while the voltmeters are documented in the Keithley 2001 Manual to express over 10 GOhm of resistance (essentially infinite resistance). The latter only occurs when the input voltage is below 20 V. The voltages at rings 181 and 182 are above 20 V and below 200 V (though actually at negative voltage), so their voltage is scaled by the shown 1.11 and 10.0 MOhm resistors to be stepped down by a factor of x10. The readings are then multiplied by x10 before being recorded in the Slow Controls database.

After Run 9, this box was disconnected and resistances were measured for the OFCW portion. Because the resistors were not separated from each other, equivalent resistances were actually measured. In the below math, R182eq refers to the resistance measured across resistor R_182, while Rfull is the resistance measured between the input to the box from ring 181 to the output for the ammeter. R111 is the combined 10 + 1.11 MOhm pair.

double R111 = 10.0 + 1./(1./1.11 + 1./10000.)
11.11
double R182eq = 0.533 // measured
double Rfull = 2.31   // measured

double pa = R111 - Rfull
double pb = R111*(R111 - 2*Rfull)
double pc = (R182eq - Rfull)*R111*R111

double R181 = (-pb + sqrt(pb*pb - 4*pa*pc))/(2*pa)
2.3614
double R182 = 1./(1./R182eq - 1./R111 - 1./(R181+R111))
0.5841

double R182b = 1./(1./R182 + 1./R111)

double Vcm = 27960.
double Rtot = 364.44   // full chain
double Inorm = Vcm/Rtot
76.720
double V181_norm = Rfull * Inorm
177.22
double V182_norm = V181_norm * R182b/(R181 + R182b)
33.724

double Rshorted = 1./(1./R111 + 1./R182 + 1./R111)
0.5286
double Rmiss = Rfull - Rshorted
1.7814
double Ishorted = Vcm/(Rtot - Rmiss)
77.097
double V181_short = Rshorted * Ishorted
40.750
double Iexcess = Ishorted - Inorm
0.377

Note that many of these numbers would be different for the inner field cages.

External resistors to make up for missing resistance can also be added to the chain here.

GridLeak Studies

Floating Grid Wire Studies

Distortions in TPC data (track residuals) are seen in 2004-2006 data which are hypothesized to come from a floating Gating Grid wire. Notes from a meeting of TPC experts regarding the topic held in October 2006 can be found here.

See PPT attachment for simulations of floating grid wires from Nikolai Smirnov which show that the data is consistent with two floating -190V wires in sector 8, and two floating -40V wires in sector 3 (all wires are at -115V when the grid is "open").

GridLeak Distortion Maps

 Using the code in StMagUtilities, these are maps of the GridLeak distortion.

First, this is a basic plot of the distortion on a series of hits going straight up the middle of a sector (black: original hits; red: distorted hits). The vertical axis is distance from the center of the TPC (local Y) [cm], and the horizontal axis is distance from the line along the center of the sector (local X). Units are not shown on the horizontal axis because the magnitude of the distortion is dependence on the GridLeak ion charge density, which is variable.

The scale of the above plot is deceptive in not showing that there is some distortion in the radial direction as well as the r-phi direction. The next pair of plots show a map of the distortion [arb. units] in the direction orthogonal to padrows (left) and along the padrows (right) versus local Y and angle from the line going up the center of the sector (local φ) [degrees].

One can see that the distortion is on the order of x2 larger along the padrows than orthogonal to the padrows. Also, it is clear that there is a small variance in magnitude of the distortion across the face of the sectors.

The next plot shows the magnitude of the distortion [arb. units] along the padrow at the middle of the sectors vs. local Y [cm] and global Z [cm]. The distortion is largest near the central membrane (Z=0) and goes to zero at the endcaps (|Z| ≅ 205 cm), with a linear Z dependence in between, which flattens off at the central membrane and endcap due to boundary conditions that the perturbative potentials are due to charge in the volume and are constrained to zero at these surfaces.

GridLeak Simulations

Nikolai Smirnov & Alexei Lebedev:
Data for STAR TPC supersector.   05.05.2005  07.11.2005
Jon Wirth, who build all sectors provide these data.

Gated Grid Wires: 0.075mm Be Cu, Au plated, spacing 1mm
Outer Sector 689 wires, Inner Sector 681 wires. Total 1370 wires per sector

Cathode Grid Wires: 0.075mm Be Cu, Au plated, spacing 1mm
Outer Sector 689 wires, Inner Sector 681 wires. Total 1370 wires per sector

Anode Grid Wires:0.020mm W, Au plated, spacing 4mm
Outer Sector 170 wires, Inner Sector 168 wires. Total 338 wires per sector
Last Anode Wires: 0.125mm Be Cu, Au plated
Outer Sector 2 wires, Inner Sector 2 wires. Total 4 wires per sector

We are most interested in the gap between Inner and Outer sector, where ion leak is important for space charge. On fig. 1 wires set is shown. The distance between inner and outer gating grid is 16.00 mm. When Grid Gate is closed, the border wires in Inner and Outer sectors have -40V, each next wire have -190V and after this pattern preserved in whole sector - see fig. 2. When gating grid is open, each wire in gating grid have the same potential -115V. Above grid plane we have a drift volume with E~134V/cm to move electrons from tracks to sectors and repulse ions to central membrane. Cathode plane has zero voltage, while anode wires for outer sector holds +1390V and for inner sector +1170V.


Fig. 1. Wire structure between Inner and Outer sector.


Fig.2 Voltages applied to Gating Grid with grid closed.

Another configurations of voltages on gating grid wires are presented on fig.3. All these voltages are possible by changing wire connections in gating grid driver. Garfield simulations should be performed for all to find a minimum ion leak.


Fig.3 Different voltages on closed gating grid (top: inverted, bottom: mixed).

This is a key for Nikolai's files: there are 4 sets of files in each set there is simulation for Gating Grid voltages on last wires. Additionally he artificially put a ground shield on the level of cathode plane and simulated collection for last-thick anode wire and also ground shield and last thin anode wire.

Setups: Standard Inverted Mixed Ground Strip Ground Strip
and Wire
Equipotentials PS

PS

PS

PS

PS

Electron paths PS

PS

PS

PS

PS

Ion paths
(inner sector)
PS

PS

PS

PS

PS

Ion paths
(outer sector)
PS

PS

PS

PS

PS

GridLeak: Gain Study

In February 2005, Blair took some special runs with altered TPC gains so we could study the effect on the GridLeak distortion. What is shown in the following plots is the distortion (represented by the profile of residuals at padrows 12 and 14, which amounts to 0.5 * [residual at padrow 12 + residual at padrow 14]) as a function of Z in the TPC. Black points are the distortions from sectors 7-24, and red are 1-6, which are the sectors where the TPC gain was reduced. I have chosen as labels "Norm", "Half", and "Low" for these three conditions of no gain change, half gain change on the inner pads only, and half gain change on both inner and outer.

First note: the ever-present problem that our model goes to zero distortion at the endcaps, while the measured distortions do not appear to do so (though the distortion curves should flatten out [as seen in the above plots] as a function of Z near the endcap and central membrane due to boundary conditions on the fields in the TPC).

Second note: I have not excluded sector 20 from these plots, which is partly to explain why the east half (z<0) has slightly less distortions than the west in these profile plots. In reality, east and west were about even for a normal run (distortions excluding sector 20).

Here is the z-phi plot for Low (it's almost difficult to see the distortion reduction in the z-phi (in "o'clock") plot for Half):

Third note: (though not too important for this study because we generally ignore east/west comparisons) the sectors between 1-6 o'clock already tend to show somewhat less distortion than the sectors at 7-12 o'clock, and because it is true on both halves of the TPC, it is more likely to be due SpaceCharge azimuthal anisotropy than asymmetries in the endplanes. Here are the distortions at |z|<50 for east (red) and west (blue) as a function of phi in "o'clock" where one can see the already present asymmetry, explaining why sectors 1-6 are already less distorted in the Norm run than sectors 7-12:

We have to normalize to sectors 7-12 to see the drop in distortions as it the runs are taken at different times when the luminosity of the machine, and therefore the distortion normalizations, are different. Here are the ratios from sectors 1-6 / sectors 7-12:

And the double ratios to see the drop in the Half and Low runs w.r.t. the Norm run:

These plots show ratios in the Z = 25-150cm range of about 0.86 and 0.59 respectively, or reductions of about 14% and 41% give or take a few percentage points. Data beyond 150cm tends to be poor and there's little reason to believe that the ratio really changes by much there. However, there does appear to be some shape to the data, which is not understood at this time.

Another way to calculate the difference in distortions is to take a linear fit to the slope of the distortions between z = 25-150cm. Those fit slopes are:

Low:
1-6: 0.000250 +/- 0.000019
7-12: 0.000420 +/- 0.000024
Half:
1-6: 0.000365 +/- 0.000023
7-12: 0.000433 +/- 0.000025
Norm:
1-6: 0.000401 +/- 0.000024
7-12: 0.000422 +/- 0.000023
Again, we need the ratio of ratios:
[Half(1-6)/Half(7-12)] / [(Norm(1-6)/Norm(7-12)] = 0.89+/-0.11 (12%)
[Low(1-6)/Low(7-12)] / [(Norm(1-6)/Norm(7-12)] = 0.63+/-0.08 (12%)
Inner reduction = (11 +/- 11)%
Outer reduction = (26 +/- 11)%
Total reduction = (37 +/- 8)%
These numbers are smaller than the reductions indicate by the above plot of double ratios of the distortions themselves. This likely reflects the errors in fitting the slopes. In that sense, the plot values may be more accurate. We need not worry in this study about getting the reduction numbers exact, but it is perhaps accurate enough to say that the inner sector gain drop reduces the distortion by about 13%, and the outer sector gain drop reduces it further by about 27% (about twice as much as the inner) from the original distortion. It is clear that both inner and outer TPC sectors contribute to the distortion, and that the outer TPC contributes significantly more to the distortion. If hardware improvements can only be implemented for either the inner or outer, then the outer is the optimal choice in this respect. It is not obvious offhand whether this is consistent with the GridLeak Simulations which we have done so far for these ion leaks.


Gene Van Buren

Long term life time and future of the STAR TPC

In an effort to better understand the future of the TPC in the high luminosity regime, several meetings were held and efforts carried through.

Present at the meeting(s) were: myself (Jerome Lauret), Alexei Lebedev, Yuri Fisyak, Jim Thomas, Howard Wiemann, Blair Stringfellow, Tonko Ljubicic, Gene van Buren, Nikolai Smirnof, Wayne Betts (If I forgot anyone, let me know).

It is noteworthy to mention that the TPC future was initially addressed at the Yale workshop in (Workshop on STAR Near Term Upgrades) where Gene van Buren gave a presentation on TPC - status of calibration, space charge studies, life time issues. While the summary was positive (new method for calibration etc...) the track density and appearance of new distorsions as the lumisoity ramped up remained a concern. As a reminder, we include here a graph of the initial Roser luminosity projection.

Roser luminosity
which was build from the following data  

200220032004200520062007 20082009201020112012 20132014
Peak Au Luminosity 4 12 16 24 32 32 32 32 32 48 65 83 83
Average Au Store Luminosity 1 3 4 6 8 8 8 8 16 32 55 80 80
Total Au ions/ring [10^10] 4 8 10 11 12 12 12 12 12 12 12 12 12

An update summary of the status of the TPC was given by Jim Thomas at our February 2005 collaboration meeting (A brief look at the future evolution of TPC track distortions ; see attachement below) to adress ongoing GridLeak issues. Slide 19 summary for teh future is added here:
Will the TPC Last Forever?

  1. Dynamic distortions driven by L
    1. 2x increase is feasible and this takes us to 2010 and (probably) beyond.
  2. Some static distortions need work
    1. e.g. Central Membrane is not flat
      Probably of academic interest
    2. Unlikely that any of these static unresolved issues will affect the useful lifetime of the TPC
  3. Beam backgrounds and ghost beams may be a problem
    1. PHENIX put up shielding
    2. Gene see’s some bad runs
  4. TPC replacement parts will eventually be a problem
  5. TPC replacement people will definitely be a problem

Questions arose as per the liability of the detector itself i.e., aging issues (including shorts and side effects) were raised along with increasing concerns related to grid leak handling. Alexei Lebedev proposed a serie of hardware modification in May 2005 to account for those issues (see What we can do with TPC while FEE are in upgrade attachement).

Relevant to possible software and hardware solutions for the grid leak are GridLeak Simulations of the fields and particle paths in the region near the inner/outer TPC sector gap.

Extensive analysis are also available from the pages SpaceCharge and GridLeak and especially (for AuAu) this page QA: High Luminosity and the Future.

Anode wire aging

 See the attached file: RD51-Note-2009-002.pdf

SpaceCharge studies

 Studies of SpaceCharge in the TPC.

Potential distortions at projected luminosities for different species

Here I show the expected distortions in the STAR TPC as represented by the pointing error of tracks to the primary vertex (otherwise known as the DCA) due to experienced (star symbols) and projected luminosities (triangle and square symbols) at RHIC. One can think of these distortions as being caused by the accumulated ionization in the TPC due to charged particles traversing it from collisions (ignoring any background contributions), or the charge loading of the TPC.

These measurements and projections are current as of November 2008 with the exception of the pp500 (pp collisions at √s = 500 GeV) projections (open symbols) [1,2]. In 2005 the expectation was that RHIC II could achieve pp200 peak luminosities of 150 x 1030 cm-2sec-1 and pp500 peak luminosities of 750 x 1030 cm-2sec-1 [3]. Presently (2008), the pp200 peak luminosity estimate at RHIC II has dropped to 70 x 1030 cm-2sec-1, and this is what is used in my plot. I have found no current estimate for pp500. The 2005 estimate indicated a factor of x5 more peak luminosity from the 500 GeV running than 200 GeV; I do not know if that is still possible, but I have chosen to use only a factor of x2 (should be conservative, right?) in the plot shown here.

I have also had to estimate the conversion of luminosity into load in the TPC for pp500 as we have not yet run this way. We did run at √s of approximately 405 GeV in June 2005, which showed a 17% increase in load over pp200 per the same BBC coincidence rate [pp400 (2005)]. I have roughly estimate that this means approximately a 25% increase in TPC load going from pp200 to pp500, with no serious basis for doing so.

Finally, the documented projections show that RHIC II will achieve approximately a factor of x2.6 increase over RHIC I for both AuAu200 and pp200 [2]. I have taken the liberty to apply the same factor of x2.6 to CuCu200, dAu200, and pp500.

(Note: The pp data use a different horizontal axis which was adjusted to lie amidst the ion data for ease of comparison, not for any physically justified reasons.)

 

 

Of the data we've taken so far, AuAu collisions have presented the highest loading of the TPC. However, CuCu collisions have the potential to introduce the highest loading for ions, and the RHIC II projections lead to pointing errors close to 20 cm! Actually, SiSi may be even worse (if we ever choose to collide it), as the luminosity is expected to achieve a factor x4.2 higher than CuCu, while the load per collision may be in the 40-50% ballpark of CuCu [2,3].

Loading due to pp collisions at 200 GeV are generally similar to AuAu collisions at 200 GeV, but 500 GeV pp collisions will load the TPC much more severely. Even in RHIC I, we may have pointing errors between 5-10 cm using my conservative estimate of x2 for the luminosity increase over pp200. If the factor of x5 is indeed possible, then use of the TPC in pp500 at RHIC II seems inconceivable to me, as the simple math used here would completely break down (the first TPC padrows are at 60cm, while 2.5*25.1 cm is larger than that)**. Even with a factor of x2, things may be quite problematic for pp500 in RHIC II.

 

References:

  1. RHIC Run Overview
  2. RHIC Projections
  3. RHIC II - Ion Operation
  4. pp400 (2005)

Some additional discussions from 2005 regarding the TPC and increased luminosities are documented at Long term life time and future of the STAR TPC.

Note: This is an update of essentially the same plot shown in the 2007 DAQ1000 workshop (see page 27 of the S&C presentation).

** Actually, a rather major point: we never get full length tracks with DCAs to the primary vertex of more than a couple cm anyhow, due to the GridLeak distortion. These tracks get split and the DCAs go crazy at that point. Since GridLeak distortion corrections are only applied when SpaceCharge corrections are, an roughly appropriate SpaceCharge+GridLeak correction must be applied first to even find reasonably good tracks from which to determine the calibration.


G. Van Buren

 

SpaceCharge in dAu

In Run VIII, dAu data was acquired at high enough luminosities to worry about SpaceCharge (it was ignored previously in the Run III dAu data (2003)). Attached is initial work by Jim Thomas to determine the charge distribution in dAu.

TPC Sector Alignment Using Particle Tracks for Run 8

TPC Sector Alignment Using Particle Tracks for Run 8

TPC group meetings 2020-

Provide link to minutes of meetings held.

Earlier minutes from during construction and commisioning can be found on https://drupal.star.bnl.gov/STAR/subsys/upgr/itpc/itpc-meetings/itpc-minutes

Started  on 10/14/2020;   but it has not been strictly kept up to date.

March 23, 2023

Present: Alexei, Flemming, Jim, Tommy, Prashanth, Yuri, Gene, and myself
 
== Hardware
— Working on lasers 
— east alignment is finished
— both lasers are working and operational
— spare pump is in hand
— checked cathode today
— no shorts so far
— finishing structure to hold magnetic coils
— next week east poletip going in (on 28th)
— gas flow starting April 18th per operations meeting
— magnet polarity switching time about 3-5 hours (related to discussion last week)
— (Prashanth) now buying gas from Lindy (sp?) 
— Prashanth ordered all needed gases
— AirGas actually supplying methane since Lindy couldn't deliver
— also ordered another 2 6-packs from different company 
— Alexei suggested another possibility, Mathison (or Madison), Prashanth will follow-up
 
== Electronics
— summary in slides from yesterday
— just one RDO remains with some fuses to change
== Software
— information from Irakli
— will provide resolution vs. momentum for new and old alignment soon?
— update on O+O 200 GeV (RFF vs FF observed difference)
— revisited RFF O+O from August-September 2022
— see plots in agenda for today's meeting
— evidence from FF for an approximately 1 mrad rotation of TPC
— looked at vertices reconstructed from tracks in east only and west only in low lumi fixed target data
— see approximately 2 time bucket shift between east and west for fixed target (as reported earlier by Xiangli)
— see additional plot in meeting agenda
— Temple University transitioning TPC work from Babu to Richard Thrutchley
— will deal with Run 22 calibration
— good positioning for this work due to previous use of calibration in calorimeter for TPC calibration
 
== AOB
— Yuri updated RDO mask and dead FEE map for all BES data sets (fixed target runs only??)
— need to check if any issues remain for '20 and '21 sets

Friday: East side was finished
from Christian

Finished up around 2:30.
-14-4 RDO was removed, replaced with working alternate
-14-2 & 14-3 RDO cables had nylon braid put in over the splices and were put back in the cable tray
 
-17-1-3 FEE and it's cable were replaced (again), still failed; moved from -3 to -5 and it worked, Tonko will update mapping
Tonko checked all of sector 14 and RDO 17-1.
 
I looked at the other cable trays and didn't see any obvious problems with the cables and water lines.
We should be ok to insert the poletip whenever.

March 16, 2023

TPC meeting March 16,2023
Present:  Gene, Tommy, Yuri, Alexei, Jim and Flemming

-Hardware
 Alexei complete East laser alignment; Laser system is now ready for upcoming run
 Next week will install ‘new’ magnetic sensors for additional monitoring.
 Q: can these be interfaced to Slow Control (Monitoring)

-    Electronics
-    The TPX s8-3 RDO issue turned out to be a network card problem in the DAQ PC. West is ok a few FEEs are masked off for run
-    East TPC should also be ready.

Gating grid tests
-    Gene is proposing that we test the effectiveness (transparency) of the GG under different beam conditions.
-    Aim to take short runs (<1 min) with 2,3,4,5kHz rate for MB data taking and similar for High lumi running later in the runs.
-    Gene will talk to Jeff on setting up run configs so it can easily be run by crew. And with JG on beam conditions.
-    The equest will be added to the TPC reafiness status

Pre run calibrations
Had long discussion on FF/RFF information that should be extracted. STAR magnet is currently wired for FF.
Yuri would like, now that Alexei has installed detector to monitor possible movement , that we switch a couple of times between FF , 0 , and RFF to see if a) there is a movement and b) if it is reproduceable. This does not need COSMIC data
It was not clear how this knowledge would translate into improved calibrations of TPC. Jim pointed out that (some) E-B mis alignment is already included in space charge distortions (static?). Gene commented that the main geometry calibration issue/uncertainties is related to the super sector to super sector alignment.
Yuri iterated that understanding if there are movements of TPC between FF and RFF will help for calibrations methods.
WWill bring the proposal up at readiness for TPC, but expect to get push back to the ‘cost’ (time, people,..) of doing multiple filed switches.

De/dx calibrations
Yuri pointed to the presentation at last S&C meetins that identified at FF/RFF difference between a correction that has been used previous for de/dx vs. position along wire attributed to wire to pad plane diff. See https://drupal.star.bnl.gov/STAR/system/files/RunXXI%20OO200GeV.pdf

FV raised the issue why this shows up in OO at 200 GeV which has about equal positive and negative charge tracks with opposite curvatures.

+ Run 21 dAu200 SpaceCharge
  - Babu continues to find essentially no asymmetry between east and west, despite historically seeing ~20% more in east
  - Gene sees asymmetry in the lower luminosity Run 21 dAu data, but it disappears at high luminosity
  - Recommendation from Gene was for Babu to process more high luminosity data to help investigate dependencies further

- Run 22 pp500 SpaceCharge
  - Babu has plenty of SpaceCharge work to do even without Run 22, plus his physics analysis
  - Gene has asked the PWGs to provide an additional helper to investigate possible fill-by-vill variations as seen in Run 17
  - Good to educate more people about SpaceCharge anyhow, even if Babu does find  time to work on it

+ New cluster finder
  - Tonko expects it to be not just different, but "better"
  - Tommy proposed a test with embedding to see if momentum resolution and acceptance improves as Tonko hopes
  - Yuri suggests not to focus on momentum resolution, but rather efficiency
  - Tommy will discuss with Tonko more

+ dE/dx adjustments by analyzers
  - Tommy brought up artificial, empirical dE/dx nsigma shifting by analyzers for BES-I datasets
  - Yuri says that's ok as a band-aid for older productions of iTPC era data, but that the new dE/dx model introduced last year should help avoid this
  - Tommy is willing to work on a standard version of the empirical band-aid for people to use, and no one objected

March 9, 2023

Present: Flemming, Yuri, Gene, Tim, and myself
 
== Hardware
— SC3 very slow (controlling Gating Grid driver)
— several things have been tried but nothing successful yet
— may ask Wayne to reboot once more
 
== Electronics
— Discussion with Christian this morning
— east end work is done except for a few cable wraps
— problem since STGC installation
— fiber connection to one TPC sector 8 RDO 3 not working
— whole RDO is out
— will be investigated
— worst case would have to reopen poletip and fix connection
— but that should be a last resort and the decision would be based on size of dead area
 
== Software
— Babu working on d+Au SC and GL for Run 21
— Yuri showed slides on TPC23 cluster finder
— see https://www.star.bnl.gov/~fiscal/star/TPC/TPC23/
— differences in number of found clusters in outer sectors between online and TPC23
— different finders putting different bounds on found clusters resulting in more from TPC23
— efficiencies are lower for TPC23 at low-pt, high eta
— result is some track losses at high eta for TPC23 compared to online but losses are small and likely acceptable
— Guanan working on TpcRs
— still looking at differences in data width compared to simulation
— Yuri will send URL to TPC list
— Yuri still needs position of TPC with respect to magnet from survey
— Also waiting on comparison of momentum resolution from old and new alignment from Irakli
— also see https://drupal.star.bnl.gov/STAR/meetings/STAR-Collaboration-Meeting-Spring-2023/Pleanary-Session-I/iTPC-performance
 
== AOB
— Flemming will give an update to Operations meeting on status of TPC on March 22nd




February 23, 2023

Present: Alexei, Flemming, Yuri, Gene, Jim, Tommy, and myself
 
== Hardware
— East laser power supply had a leak
— pump was leaking
— have a spare and beginning repair today
— TPC water system has leak from spiral hose as well
— was repaired yesterday
— notified electronics folks that they may resume work
— some discussion about how to align with movement of TPC
 
== Electronics
— Nothing new reported this week
 
== Software
— Babu working on d+Au SC and GL for Run 21
— Run 21 O+O large calibration production is finishing today
— can provide to Yuri to look at dE/dx
— For O+O Yuri still has a puzzle from survey
— difference between FF and RFF
— factor 2 deviation was reduced but some residual still present
— looking at 2011 27 GeV FF data to see if rotation of TPC is there 
— TPC23 cluster finder 
— using same DAQ file for online and offline
— Yuri sees some difference in TPX outer sector(s)
— comparison of online and (TPC23) offline shows about 10% difference in clusters
— checking to see if difference is due to gain file
— Yuri says Irakli reported to have finished alignment with his new approach
— report is there is a big improvement in alignment, but need to hear from Irakli directly
— Yuri needs position of TPC with respect to magnet from survey
— Guanan has finished tune-up of TpcRs
— still some question about tails
— data slightly wider than simulation in time direction
— want to understand that
— next will be to check data for fixed target runs
 
== AOB
— Some discussion of plan for fixed target (with 3 new targets)
— we would like to see the drawings of the location(s) of new targets on east side
— Tonko back in Croatia
— software expected to allow us to start with ~4 kHz
— will get to 5 kHz a little later


February 16, 2023

Present: Yuri, Gene, Tommy, and myself (Richard)
 
== Hardware
— Nothing new reported this week
 
== Electronics
— Nothing new reported this week
 
== Software
— Babu working on O+O and d+Au SC and GL
— FF redone due to issue with tables
— high luminosity looks good
— low luimi shows some discrepencies
— will revisit once alignment is done
— can start large calibration production for O+O as early as tomorrow
— d+Au needs more attention once 
— Yuri has some questions for Alexei on survey
— Guanan working on TpcRs
— investigating some issue with prompt hits
— Noticed Tonko has updated cluster finder
— some mismatch between inner and outer sectors
— misses last two padrows (71 and 72)
— believe there is problem with pad maps



February 9, 2023

Feb 9, 2023 TPC meeting

Present: Yuri, Prahsanth, Alexei, Flemming, Tommy

--hardware
Westpole installed. #d sensor installed before and can be viewed so we can observe movement when filed is turned on, changed polarity etc.
Water leak in East laser box.  Alexei will investigate.

-    After insertion 2 bad FFEs identified. Likely due to squeezed cable. Need to ensure better pre-closure inspection for future years.

Software
  Tommy and Tonko identified one source of differences between online and offline clusters. Namely simulation of hardware asic that merged time sequences separated by just one empty timebin (below threshold) would be merged into one

Yuri reported that analysis of recent survey revealed that TPC is rotated relative to RFF.
Still needs to talk to survey group to understand coordinate systems.

Yuri reported (see meeting page) that the QA plots for pbar vs p vs pt for the period c (production nomenclature) ie. From after covid shutdown is different than before shutdown, Period “” and b .

January 25, 2023

Present: Alexei, Yuri, Gene, Flemming, Tommy, Prashanth, and myself (Richard)
 
== Hardware
— 3D sensor installed on west side 
— glue is now dry, need to install wire
— expect a couple hundred microns accuracy
— poletip to be install Monday or Tuesday next week
— now shorts detected so far
 
 
== Electronics
— Tonko is done with checking west side electronics
— waiting for poletip installation for final check 
— if check is okay will wait for run
— testing will occur at regular intervals 
— all RDOs on east have been checked
— some FEEs remain to be checked and possibly replaced
— Tonko working on PC programming for DAQ speed-up work
 
== Software
— Gene found a workaround for SC and GL calibration jobs
— Babu paused dAu work to work on O+O
— RFF and FF OO calibrated as of Tuesday evening
— large calibration production will be ready to go shortly (once values are in dB)
— Richard will reach out to Irakli to ask about alignment work (CC Gene)
— Yuri working on understanding differences between data and simulation timing tails
— simulation tails are longer than in data
— using 2009 data (with no shaper)
— Also working on data from survey group
— Yuri may be out for a couple weeks 
— Gene will be out week after next
— Tommy presented work on discrepancies between online and offline clusters
— shows comparison of clusters found by online and offline
— seems online CF is merging some clusters that offline finds separately
— one possibility mentioned is that dead pad mask may make a difference in how potential clusters are connected
 

January 11, 2023

Present: Alexei, Yuri, Mike, Gene, Flemming, Tommy, and myself
 
== Hardware
— survey finished with new pins
— will begin laser work once current measurements are done
— Tonko has been working with electronics since Monday
— still have not received documents from survey group
— will reach out to see what's happening with documents
 
 
== Electronics
— Tonko is presently at BNL
— going through electronics on both west and east ends
— just a couple RDOs have some problems
— hoping to have that work done by end of week
 
== Software
— Babu still facing some issues with SC and GL work
— O+O will still be higher priority than d+Au, so Babu will focus on that next
— Yuri says TpcRs tuning continues
— need to understand relation between actual chip parameters and those in the dB
— some question regarding signal shape on anodes wires and ALTRO chip tail cancellation
— Tommy working on online and offline cluster finding match
— trying to find mismatched clusters and recover pixel information
— difficult work, at text file level
— stay tuned




January 4, 2023

Present: Alexei, Yuri, Prashanth, Tommy, Gene, Flemming, and myself
 
== Hardware
— east side is finished from bake-out
— will start now fixing RDOs on east side
— also seems that survey is done on east side but Alexei has not received official word yet
 
== Electronics
— some delay in work waiting on beam pipe bake-out heat to dissipate 
 
== Software
— Babu still facing some issues with SC and GL work
— O+O will still be higher priority than d+Au, so Babu will focus on that next
— final SC work should use final version of alignment from Irakli
— but should have the interim alignment results should 
— 9.2 and 11.5 GeV dE/dx calibration finished
— checked with standard library, no diffs with dev
— dB timestamp should be taken from this year
— link to current status of work with new model (dN/dx)
— see https://drupal.star.bnl.gov/STAR/system/files/Revison%20STAR%20TPC%20dEdx%20Model_0.pdf
                
— see https://drupal.star.bnl.gov/STAR/system/files/Revison%20STAR%20TPC%20dEdx%20Model_0.pdf







December 21, 2022

Present: Alexei, Yuri, Guannan, Tommy, Gene, Flemming, Irakli, and myself

== Hardware
    — survey done on west but not east side
        — Alexei (currently on vacation) trying to push a little to get it done
        — found good way to put cameras for TPC movement
        — Alexei will prepare short presentation upon return == Electronics
    — inner sector RDOs on west side have been put backup by Alexei
    — but on east side some RDOs still need to be returned to proper position         — Alexei needs to talk to Tonko about east side before Tonko resumes work with RDOs on or after 1/1/23

== Software
    — Gene mentions that Babu is still facing some issues on SC and GL
    — PWGs now prioritizing O+O
        — SC is ongoing for that but Babu needs some help from Gene
        — once that’s done will move forward with TPC dE/dx etc.
        — Irakli working on alignment, RFF seems to be converged
            — FF does not seem to be in right place in X after RFF convergence (within 100 um)
    — Yuri finished with dE/dx calibration (7.7 and 9.2 from 2020)
        — have put calibration constants from new model in dB
        — need final check with Standard Library
    — Guannan is continuing work on deep tuning of TpcRS
        — some problems observed
    — Tommy continuing with offline/online cluster comparison
        — 0.5% difference observed and must be understood

December 14, 2022
No meeting, but Yuri presented an update on dE/dx at the s&c meeting  - presentation
Electronics status presented by Tonko at analysis meeting - presentation

December 7, 2022

Present: Alexei, Yuri, Gene, Flemming, Mike, and myself
 
== Hardware
— finished cleaning of all TPC except 12 points on east side
— interrupted by water leak
— expected that leak will be fully repaired this week
— Tonko nothing with TPC until Jan. 1st
— plenty of time for Alexei to finish work
 
== Electronics
— nothing new reported this week
 
== Software
— Yuri finished with dE/dx calibration (7.7 and 9.2 from 2020)
— will put calibration constant in dB and request production
— will prepare calibration for entire BES data set
— some other activities ongoing around combined PID (from all detectors)
— Guannan is continuing work on deep tuning of TpcRS
— Gene mentioned Babu has started working on SC and GL
 
 
== AOB
— Flemming mentioned minutes from QA board
— Run 20192001 shows anomalous drop in sector 1 in phi vs. eta
— not easily seen in QA plots (small)
— sector 1 had many auto-recoveries in that run
— seems that one RDO did not come back fully resulting in an inefficiency
— likely need a new QA plot to help identify this problem in the future


November 30,2022

Present: Alexei, Yuri, Irakli, Gene, Tommy, Guannan, Flemming, and myself
 
== Hardware
— Tonko finished yesterday
— much easier for Alexei to work on holes now
— expects to finish cleaning today or tomorrow
— needs to invite survey group to make extensive plan
— moving scaffolds, etc.
— need to make 3D tool to move fiber
— Flemming asked about water leak
— exact location has yet to be identified
— Robert Pak was notified (instead of Alexei?)
— CAD requested some parts to fix leak properly
— Alexei finished cleaning but would like to work on whole inner sector 
— last 2 RDOs will shift a little bit (approximately 0.5 inch)
 
== Electronics
— talked to Christian on Thursday, some issues remain
— Tonko switched off all electronics yesterday
— however, this morning sector 15 (or 16 or 17 was found to be on)
 
== Software
— Irakli presents alignment work on '21 cosmics data
— previously kept outer sectors fixed and aligns inner sectors to them
— some issues found so now reversed procedure (fix inner and align outer)
— related to lack of knowledge about exact position of pad plane on outer sector
— we have a good survey of pad plane position on sector for inner (iTPC) sectors 
— so, inner sectors are now being used as reference instead of outer
— deltaZ and deltaY converge quickly, deltaX is slower but seems to be converging with continued iterations
— iterations will continue
— supersectors still to come
— Gene on production priorities
— Run20 11.5 and 9.2, then 200GeV d+Au from Run21(SC and GL need to be finished)
— next comes 200 GeV O+O (has 2 field settings)
— need to understand static distortion and alignment issues  beforehand
— working groups want a test production anyway
— will finish SC and GL for O+O
— then will move forward with preview production with whatever corrections state exists at the time
— Tommy working on automatic pull of histograms for QA Fast Offline analysis
— finishing work on differences in cluster finding results in simulation and production
— Yuri continuing iteration on constant dE/dx work (each needs a couple days)
— Guannan work on tuning TpcRS
— some discrepancy in clusters noticed by Xianglei 
— needs to be understood
 

November 16, 2022

Present: Alexei, Yuri, Tommy, Geydar, Vinh, Gene, and myself
 
== Hardware
— checked almost all holes
— need to coordinate with Tonko now about non-interference with electronics work
— did not get to check 3D sensor fit yet due to activity on west side
 
== Electronics
— nothing new for this week
 
== Software
— presentation from Vinh on dE/dx QA from 7.7 GeV
— see https://drupal.star.bnl.gov/STAR/blog/lbavinh/dEdx-QA-AuAu-77-GeV-Run21
— conclusion is that ToF m^2 cuts give, at best, only factor 10 improvement in reduction of misidentification (see talk)
— Gene asked if ToF group has studied the effect of tails on the distributions
— ultimately attempting to do combined PID with all detectors
— this is first step comparing TPC and ToF
— Yuri has frozen new modified dE/dx model
— calculating dE/dx from all BES data
— continuing iteration and work on methodology
— expect new dE/dx calibration for all BES II data soon, but not clear how soon
— Gene asked about choice of TFG22e library vs. SL22b
— only TFG offers KFParticle finder



November 09,2022

Present: Flemming, Alexei, Tommy, Mike, Gene, and myself
 
== Hardware
— continuing to clean TPC wheels for survey
— about 70% done
— Tonko is working inner sector electronics, so being careful not to interfere
— 3D sensor prototype has been built, looks okay
— want to put on TPC to check fit
— if okay will request manufacture of a support structure
 
== Electronics
— talked with Christian
— Christian has list of FEE cards to look at (about 10 in iTPC)
— few will need to be changed
— seems like mostly cable issues
— everything still moving forward 
— Alexei needs some access to some areas being used 
— need access once for testing and then again for mounting
— need to coordinate with Tonko to deconflict access
— will first finish with holes to which he currently has access
 
 
== Software
— reached out for space charge work 
— project with Rice students on PID
— students should join TPC group for discussions
— Richard will check with Yuri to ensure the right names are added to the TPC list
— Gene mentioned the code updates to cluster finder from Tommy and Yuri look good
— some tuning still to be done

November 02, 2022
present:
Gene, Yuri, Tommy, Alexei, Chenlian Jin, Flemming
excused: Richard

Alexei; working on probes for survey.

Open questions cosmics for survey field A/B before run?

Gene: There is no progress on space charge for some BesII years: dAu, 17.6 GeV FXT
Also need for run 22
Will contact Temple.

Nhits issues solved. Will pay more attention to cluster status when selecting/counting cluster distributions

Yuri:
a) will wrap up de/dx model for higher Z. Couple of weeks.
Request that we have a presentations for model before presenting to collaboration.

b) There is interest in making common PID (de/dx,btof,etof,...) for BESII data set from Rice group. Will like to use tpc meetings and e-mail list for this activity.
Suggestions add and invite this group

October 26, 2022

Present: Flemming, Alexei, Tonko, Yuri, Tommy, Mike, Gene, and myself
 
== Hardware
— continuing to clean holes for survey
— building 3D sensor for TPC displacement measurement
— no measurement possible until March or April due to construction
 
== Electronics
— see slide from Tonko
— expectation is everything will be done, including debugging and verifying, by March, 2023
— includes rewrite of FPGA firmware, rewrite of DAQ framework, and new cluster finder
— note that verification of TPX will require tracking (deconvoluting stage)
— should be done within next 2 weeks 
— cannot spend time on discrepancies in embedding right now, will be revisited later
— hardware should be done earlier, by end of January
— current hardware work (RDO repair, etc.) is proceeding slowly
— requests will be made regarding how to adjust the scheduling
 
== Software
— Yuri and Tommy investigating embedding
— checks have been done for Runs 18 and 19
— software commits have been made
— some issue with backward compatibility
— one line of code should fix
— similar problem to be addressed in Run 17
— work on dE/dx model continues
— Irakli continuing alignment work for Run 21

October 19,2022

Present: Flemming, Alexei, Yuri, Tommy, Mike, and myself
 
== Hardware
— determined 3D sensor to be used
— tools ordered to make extensions for CCD cameras
— need to order better scales (0.5 mm scales)
— working with survey targets
— will move some RDOs to have 4 points visible
— question to Yuri if 150 um accuracy is sufficient
 
== Electronics
— talked to Christian
— some small problems remain with the RDOs
— have water back now and some things taken to lab
— Tonko was notified that water has been restored
 
== Software
— Yuri and Tommy investigating embedding
— efficiency online and offline are different in Run 18
— due to gain table
— interface changes were made to fix loading of gain file
  — some differences remain, investigation continuing
— need 1-2 more weeks
— work on dE/dx model continues
— calibration files generated for Run 20
— Irakli joined alignment work for Run 21
— results will be forthcoming
 

October 5, 2022

Present: Flemming, Alexei, Yuri, Gene, Tommy, and myself
 
== Hardware
— still working on two options (sets of readouts)
— these are for the 3D sensor
— need a non-magnetic 3D sensor but almost no options for that
— Bill is done with the survey device
— tested and has ~50 um accuracy
— various holes around TPC need to be cleaned (filled with corrosion or debris)
— work also proceeding on TPC electronics (power supply checks, etc)
 
== Electronics
— nothing new reported this week
 
== Software
— Two issues discussed
— embedding for Run18 was first
— offline cluster finder finds more clusters than online cluster finder
— was due to gain table
— issue has returned, Tommy is working on it
— Flemming agreed the problem looks to be related to the gain table based on plots shown in the software and computing meeting
— other issue is 7.7 GeV dE/dx 
— Rice group has found some deviation
— shift is thought to be due to muons from pion decays 
— Yuri working to improve dE/dx model
— need to use data to correct model
— ultimately would want to use integrated detector subsystem information to correct model
— Rice may have one or two more student that can contribute to this effort
— also 1 more student from Dubna who will work on PID issues



 August 17, 2022

Present: Flemming, Alexei, Yuri, Mike, Tommy, and myself (Richard)
 
== Hardware
— methane leak is now something like 2 cc/min (very small) 
— Alexei can now access scaffolds for lasers
— also need to access areas that are outside scaffold volume (for 8 and 10 o'clock laser boxes)
— will need an additional platform of some kind
— talked with Bill Strubele about repeating survey with new tool
— will report back any issues
 
== Electronics 
— Tonko was waiting on feedback regarding energizing the electronics
— this is for remote operation of FEEs
— Tim will contact Tonko
— important for inner TPC sectors where replacements have been done
— Tonko believes he will be done with most work in programming FPGAs in about a month 
— can give status update at collaboration meeting
 
== Software
— discussion of some disagreement on nSigma in 7.7 GeV ('21) at Software Group meeting
— some discrepancy with regard to 
— in work on dE/dx for composites see ~3% shift between positives and negatives
— positives slightly higher than negatives
— same effect is also observed _in_simulations_ (puzzling)
 

Hardware update,
More on de/dx for 7.7 GeV auau
update from Alexei of TPC electronics readiness

About 2-3 weeks ago we had power dip and since I not paid attention to platform. I found some TPC VME crates and TPC FEE VME were OFF. I powered FEEs from platform(all PS have green LEDs) and control room, everything looks OK, but red LEDs on TPX PCs:   2, 24, 25-36 remained red.

August 10, 2022

minutes

July 27, 2022
Present: Alexei, Yuri, Gene, Tommy, and myself

== Software
 — some slides on current state of dE/dx model work
   — will be updated as work progresses
   — will post slides with current status (keeping in mind there is more to do)
  -- see https://drupal.star.bnl.gov/STAR/system/files/Revison%20STAR%20TPC%20dEdx%20Model.pdf

July 7, 2022
Hardware update. Survey done, Short repaired,
Software de/dx modle update
minutes

June 29, 2022
hardware update. Prepare for survey. Concern for Ar availability
software Tommy,Yuri have been comparing cluster finders offline/online
minutes 

June 22, 2022

Hardware update, survey. TPC electronic speedupgrade, repair  status
software update
minutes

June 15, 2022
Updates on survey plans
minutes

June 8, 2022
Updates on old vs. new cluster finder, simulated charge distribution from fixed target
minutes

June 1, 2022
Updates on dE/dx work, old vs. new cluster finder, simulated charge distribution from fixed target
minutes

May 18, 2022
TPC survey delay, inner sector wires, electronics update, 7.7 dE/dx work
minutes

May 11, 2022
gap measurments, 3D sensor prep work, inner sector wires, electronics update, 7.7 dE/dx work
minutes

April 27, 2022
nitrogen switchover, 3D sensor prep work, spare inner sectors, space charge corr., 7.7 dE/dx work
minutes
 
April 20, 2022
survey results, electronics status end of run, 7.7 de/dx calib done
minutes

April 7, 2022
prep for FF/RFF survey, space charge corr, updated manuals
minutes

February 2, 2022
discussion on added CF4 to TPC, methane shortage, run19 FEE status tables complete
minutes

January 12, 2022
to from run22 submitted, many RDO failures, look at skipping alt row for reco (pp500)
minutes

December 1, 2021
Analysis of shorts, laser response, tpc blower fixed
minutes 

November 24, 2021
Some results from short measurements, Open issues ahead of run, missing Vdrift for a day.
minutes

November 17, 2021
Magnetic field settings reading GG slow control layout, short discussion, variable vdriftd
minutes 

November 10, 2021
Plan for measurements for field cage short, 
minutes

November 3, 2021
West field cage short; run 19 space charge; upcoming calibration issues.
minutes

October 27, 2021
TPC electronics status (Tonko), Laser West repaired, 19.2 GeV space charge nearly complete.GG slow control.
minutes

October 20,2021
hardware status, GG driver, calibration update(Yuri) see slide
minutes

October 13, 2021
hardware status, cosmic request for run22 startup
minutes

September 22,2021
9.2 TPC space charge cal completed,
Embedding issue for FXT Xianglei discussed. 
minutes

September

September 8, 2021
brief meeting ; hardware cathode control
minutes

September 1, 2021
minutes

August 25, 2021 
NMR status, laser progress
de/dx short term time dependence, helpers for calibration
minutes

August 18  no meeting

August 11, 2021meeting link
agenda: hardware, calibration and software update

minutes

 August 4, 2021 meeting link
agenda:
- hardware status
-- gating grid
- calibrations
- TPC tasks for calibrations, software maintenance 
-- presentations on meeting link

minutes
 

December 9, 2020
Agenda:
-hardware status
- software
--- alignment progress 
--- https://drupal.star.bnl.gov/STAR/system/files/he34embedding4th.pdf
--- AOB

meeting link: https://drupal.star.bnl.gov/STAR/event/2020/12/09/tpc-weekly

December 2, 2020
agenda hardware status
meeting link and minutes : https://drupal.star.bnl.gov/STAR/event/2020/12/09/tpc-weekly

November 18, 2020
agenda:
hardware - SF4 tests; laser system
software; 3He efficeincies, track splitting - see talk on meeting event page
aignment progress

meeting place: https://drupal.star.bnl.gov/STAR/event/2020/11/18/tpc-weekly

November 6, 2020
Discussion points:
TPC power cables
Installation of Gating Grid reinstallation
Presentation on 3He efficiencies
Clusterfinder offline vs online
meeting place  https://drupal.star.bnl.gov/STAR/event/2020/11/04/tpc-weekly

October 28, 2020
Discussion points:
hardware: TPC gas investigation; electronics repair
software: GG turnon effects 
meeting place  https://drupal.star.bnl.gov/STAR/event/2020/10/28/tpc-weekly

October 21, 2020
Discussion points:
hardware: gating grid driver
software: Super sector alignment procedure
meeting page https://drupal.star.bnl.gov/STAR/event/2020/10/21/tpc-weekly

October  14, 2020
Discussion points
Hardware: TPX sector 13,14 powersupply cable repairs
Software:  Super sector alignment procedures 

meeting page https://drupal.star.bnl.gov/STAR/event/2020/10/14/tpc-meeting
minutes available on that page


TPC performance

 Documentation on performance of the TPC
 

 

 

TPC Hit Errors (2008)

NOTE: this was for 2008

 These are the TPC hit errors as parameterized by Victor Perevoztchikov in order to normalize pulls (and thus chi squares) of ITTF tracks versus z, dip angle, and crossing angles:

error_xy = sqrt([0]+[1]*((200.-y)/100.)/(cos(x)*cos(x))+[2]*tan(x)*tan(x))
error_z  = sqrt([0]+[1]*((200.-y)/100.)*(1.+tan(x)*tan(x))+[2]*tan(x)*tan(x))

where y in the equations is the z coordinate in the TPC, and x in the equations is the crossing angle for error_xy, and the dip angle for error_z. The parameters are:

Inner TPC:

error_xy->SetParameters(0.0004,0.0011513,0.01763);
error_z ->SetParameters(0.00093415,0.0051781,0.014985);

Outer TPC:

error_xy->SetParameters(0.0012036,0.0011156,0.063435);
error_z ->SetParameters(0.0026171,0.0045678,0.052361);

Plotting these as a function of z and crossing/dip angles gives:

Inner TPC:

Outer TPC:

EPS versions of these images are attached.

 

TPC Hit Errors (Run 9)

 As was TPC Hit Errors (2008), hit errors were found for Run 9 by Victor Perevoztchikov.

I created a macro which can be used by anyone to generate these plots, obtainable by CVS checkout from offline/users/genevb/hitErrors.C. Attached to this page are text files used as input for pp500 and pp200 (Victor ran on data with timestamps of 20090326.082803 and 20090425.093105 respectively), and eps versions of these plots.

 

pp500:

__________________________________

pp200:

 

-Gene Van Buren

TPC Pt and DCA resolution

[Update on 2010-06-21: global track momentum resolution has now also been studied using Cosmic rays study, version 2, with some as yet unexplained good performance at very high pT.]

_____________ 

Jim Thomas has written code to model transverse momentum (pT) and pointing resolution near the primary vertex (DCA) of tracks using various detectors. The code has been tuned to match TPC performance under low luminosity conditions, but assumes (for simplicity) TPC hit errors of 0.06 cm in rφ and 0.15 cm in z (which can be compared with the TPC Hit Errors (2008)), throwing tracks at η = 0.5 and including hits in "nearly all" 45 padrows (padrows 1 and 13 are dropped).

Shown below are the pt resolutions for various options:

  • Black filled triangles: embedded antiprotons at half field
  • Black open triangles: embedded pions at half field
  • Cyan: model prediction for global (anti)proton tracks at half field
  • Green: model prediction for global pion tracks at half field
  • Blue: model prediction for global pion tracks at full field
  • Red: model prediction for primary pion tracks at full field (primary vertex known to 1 mm)
  • Magenta: model prediction for primary pion tracks at full field (primary vertex known to 0.3 mm)

 

Note that the embedding data was shown in Figure 10 of the TPC NIM paper [1]. At low pT, the resolution is dominated by multiple coulomb scattering (MCS) effects. At high pT, the momentum resolution approaches a C * pT2 dependence (note that C can be considered the inverse transverse momentum resolution, as δ(1/pT) = δ(pT) / pT2 for small δ(pT)). For the above curves, C (in units of inverse momentum [c/GeV]) is approximately:

  • Cyan and Green: 0.018
  • Blue: 0.009
  • Red: 0.005
  • Magenta: 0.003

Here is the same model run with only using every other padrow of the TPC:

 For the above "less hits" curves, C is approximately 50% higher in all cases:

  • Cyan and Green: 0.027
  • Blue: 0.013
  • Red: 0.008
  • Magenta: 0.0045

Additional studies of the momentum resolution come from Yuri Fisyak for Monte Carlo simulations (similar to a thorough embedding study done using TPT in 2002 by Jen Klay) in the following plots for globals and primaries in AuAu200 and pp500 [with pileup] collisions. The comparison is best made with the blue [AuAu globals] and magenta [AuAu primaries] lines above. The data matches reasonably well with the "less hits" model curves, with the exception of an additional offset of about 0.5% for global tracks. The pileup in the pp500 simulation probably causes its results for primaries to have even further degraded resolutions than the red [pp primaries] lines above.

 

AuAu200 [Run 10 FF setup]:

 

pp500 [Run 9 RFF setup, with pileup]:


Shown below are the DCA resolutions from the same model described earlier, using tracks with "nearly all" hits, for 

  • Cyan: model prediction for global (anti)proton tracks at half field
  • Green: model prediction for global pion tracks at half field
  • Blue: model prediction for global pion tracks at full field

 

[note: I do not know what the dashed lines are]

As can be seen, the field makes only a small impact on the DCA resolution. When using "good" (high quality) tracks for calibrations, I regularly find a mean DCA resolution for global tracks in full field of just under 3 mm.

EPS versions of these images are attached.

 

References:

1. M. Anderson et al., Nucl. Instr. and Meth. A499 (2003) 659-678

TPC speed upgrade 2022

 This page is setup to keep track of meetings documents for the task force work for the
tpc speed upgrade for runs 23-25.



Status as of December 14 22,
Presentation from Tonko at collaboration meeting (pdf)

TPC meeting June 22

Update from Tonko on progress (pdf)


Meeting March 24

Discussion points send out ahead of meeting (Flemming) pdf
Progress slides - tonko  pdf
minutes - Richard pdf




First meeting February 17

overview presentation and discussion (ppt)
minutes (pdf)
-
after meeting Jeff posted the projected resource needs for handling data to S&C (pptx)


Charge and Committee

 We would like to form the STAR TPC DAQ improvement task force. With the firmware changes on the TPC electronics, it is possible for us to double the TPC electronics readout rate with a minimal cost. This will greatly enhance STAR physics capability. Our past Beam Use Request has taken the planned upgrade into account.  The long shut down anticipated after Run 22 provides us an opportunity to do the firmware change and evaluate the impact for physics data analysis. The task force is charged to evaluate the readiness of the TPC DAQ improvement for Run 23 and beyond:

 

  1. What are the resources required to realize DAQ improvement?
  2. Where is the bottleneck for this upgrade? What are the risks?
  3. What software changes (online and offline) are required to accommodate the upgrade?
  4. What hardware and network changes are required to handle this upgrade?
  5. Evaluate the impact of proposed changes on physics capabilities.
  6. What is the timeline and path toward completion of this upgrade? 
  7. Report to management regularly and provide input to beam use request for Run 23 and beyond.

 

The members are:

Flemming Videbaek (co-Chair)

Richard Witt (co-Chair)

Zhenyu Chen

Xin Dong

Yuri Fisyak

Carl Gagliardi

Jeff Landgraf

Tonko Ljubicic

Chun Yuen Tsang

Gene Van Buren

 

Lijuan and Helen

 

TPX high rate test